OpenStudy (anonymous):

the polynomial 9x^2+6X+1 has the factorization A.(9x-1)(x-1) B.9(x-3)(x+1) C.(3x+1)^2 D.the polynomial 9x^2+6x+1 cannot be factored

OpenStudy (texaschic101):

(3x + 1)^2 = (3x + 1)(3x + 1) = 9x^2 + 3x + 3x + 1 -- combine like terms 9x^2 + 6x + 1

OpenStudy (anonymous):

I think its c

OpenStudy (texaschic101):

it is

OpenStudy (anonymous):

The formula (x-y)^3=x^3-3x^2y+3xy^2-y^2 holds A.True B.false

OpenStudy (anonymous):

@texaschic101